can someone explain ( B )to me? we multiply by 100 which have only one significant number, thus the answer must have only one significant number but my friend say 0.24?3. Perform the following mathematical operations, and express each result to the correct number of significant figures a. 4.1
Answer #1

B.

8.925 - 8.904 = 0.021

Now, 0.021/8.925 = 0.0024

Hence, 0.0024*100 = 0.24

Here, 100 = 1.0*102, so it has two significant figures (1,0)

Hence, you need to express the answer with two significant figures (2,4)